OpenStudy (anonymous):

ok ... i neeed some help .. will someone please help me with ..... ; (g+4)-3g=1+g

OpenStudy (saifoo.khan):

g + 4 - 3g = 1 + g Add like terms. 4 - 2g = 1 + g -2g - g = 1 - 4 -3g = -3 3g = 3 g = 3/3 g = 1

OpenStudy (anonymous):

thanks .. what about ....... 2r-(5-r)=13+2r

OpenStudy (saifoo.khan):

2r -5 + r = 13 + 2r Solve.

OpenStudy (anonymous):

could you help me with ..... 5g+4(-5+3g)=1-g